OpenStudy (nevermind_justschool):

I will medal, thank you in advance. :) Dan had $160 in his bank account. After depositing his paycheck, he had $224 in his bank account. What was the percent of increase in Dan’s bank account after the deposit?

OpenStudy (k_lynn):

A deposit is when you put money INTO your bank account. so what do you think you need to do @nevermind_justschool

OpenStudy (jdoe0001):

well.. if we take the original amount of $160 as the 100% it went up to 224 so the difference is 224 - 160 = 64 so what percent is 64 of 160 then? \(\large \begin{array}{ccllll} amount&\% \\\hline\\ 160&100\\ 64&x \end{array}\implies \cfrac{160}{64}=\cfrac{100}{x} \) find "x"

OpenStudy (k_lynn):

Right, that's one way of doing it.

OpenStudy (nevermind_justschool):

To be honest, I am lost...

OpenStudy (k_lynn):

Wait. I read the question wrong. Sorry :/ Maybe @sammixboo , @AlexandervonHumboldt2 or @Directrix can help you.

OpenStudy (alexandervonhumboldt2):

160*(1+x/100)=224

OpenStudy (nevermind_justschool):

Oh, thank you

OpenStudy (alexandervonhumboldt2):

x is the number of %

sammixboo (sammixboo):

Well, you have